OpenStudy (gabylovesyou):

Which of the following is not a way to represent the solution of the inequality 7 - 9x - (x + 12) /= -3 x

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

What do you get when you solve for x?

OpenStudy (gabylovesyou):

-10x - 5 </= 25

OpenStudy (gabylovesyou):

right?? im not sure ^^

jimthompson5910 (jim_thompson5910):

good, keep going

OpenStudy (gabylovesyou):

mmmm x = -1/2

jimthompson5910 (jim_thompson5910):

no, add 5 to both sides

OpenStudy (anonymous):

add the 5 and then divide to get x.

jimthompson5910 (jim_thompson5910):

-10x - 5 <= 25 -10x - 5+5 <= 25 +5 -10x <= 30 keep going

OpenStudy (gabylovesyou):

its x grater than or equal to -3

OpenStudy (gabylovesyou):

x >/= -3

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

you can flip that to say -3 <= x so A and C are out

jimthompson5910 (jim_thompson5910):

keep in mind you're looking for something that is NOT equivalent (it's backwards and confusing...i know)

OpenStudy (gabylovesyou):

so would it be B?

jimthompson5910 (jim_thompson5910):

it is

jimthompson5910 (jim_thompson5910):

x <= -3 and x >= -3 are two completely different things

OpenStudy (gabylovesyou):

it is as in its not or what??? lol its confusing

jimthompson5910 (jim_thompson5910):

it's the answer because it's not equivalent...yeah tell me about it

OpenStudy (gabylovesyou):

lol ok .... :) thanks

jimthompson5910 (jim_thompson5910):

yw